### v2

# Mixed Membership Estimation for Weighted Networks

Abstract
Community detection in overlapping un-weighted networks in which nodes can belong to multiple communities is one of the most popular topics in modern network science during the last decade. However, community detection in overlapping weighted networks in which edge weights can be any real values remains a challenge. In this article, to model overlapping weighted networks with latent community memberships, we propose a generative model called the degree-corrected mixed membership distribution-free model which can be viewed as generalizing several previous models. First, we address the community membership estimation of the proposed model by an application of a spectral algorithm and establish a theoretical guarantee of consistency. We then propose overlapping weighted modularity to measure the quality of overlapping community detection for weighted networks with positive and negative edge weights. To determine the number of communities for weighted networks, we incorporate the algorithm into the overlapping weighted modularity. We demonstrate the advantages of degree-corrected mixed membership distribution-free model and overlapping weighted modularity with applications to simulated data and eleven real-world networks.
keywords:
Overlapping community detection, overlapping weighted modularity, overlapping weighted networks
[…]
[label]organization=School of Mathematics, China University of Mining and Technology,city=Xuzhou, postcode=221116, state=Jiangsu, country=China
1 Introduction
Community detection is a popular data mining problem in network science and it is to partition a network into several components, where each component is called a community (also referred to as a module or a cluster, or a block). Generally, nodes within the same community have more interactions than nodes across communities. Community detection has many real-world applications in various domains, like sociology, biology, marketing, and computer science girvan2002community ; newman2003the ; barabasi2004network ; guimera2005functional ; palla2007quantifying ; andersen2012overlapping ; garcia2018applications . For example, community detection in biological networks helps researchers to analyze the interaction of brain regions and have a better understanding of the influence of the interaction on brain functions garcia2018applications ; huang2019ocomm ; tooley2022functional . Community detection in social networks can help to identify crime organizations magalingam2015using ; legramanti2022extended . Community detection in online social networks and food social networks can facilitate recommender systems sahebi2011community ; gasparetti2021community ; rostami2023novel . Community detection in coauthorship and citation networks can be used to identify meaningful communities for scientists Newman2004 ; ji2022co .
For its wide applications, in past decades, substantial methods have been developed to learn community structure for non-overlapping un-weighted networks in which each node belongs to a single community and the edges or links between nodes are either present or not fortunato2010community ; fortunato2016community ; papadopoulos2012community , where un-weighted networks are also referred to as binary networks lu2014algorithms . For community detection in non-overlapping un-weighted networks, perhaps the most popular and well-studied generative model is the Stochastic Blockmodel (SBM) SBM . The SBM assumes that the probability of generating an edge between two nodes depends on their belonging communities. Under SBM, nodes within the same community have the same expected degree while node degrees always vary in real-world networks because there are “hub” nodes with many more edges than other nodes in the same community. To solve this limitation, the Degree-Corrected Stochastic Blockmodel (DCSBM) DCSBM was proposed by introducing node heterogeneity parameters. Many works have been developed to study SBM and DCSBM, to name a few, rohe2011spectral ; lei2015consistency ; SCORE ; joseph2016impact ; abbe2016exact ; chen2018convexified . One limitation of SBM and DCSBM is, they assume that the network has non-overlapping properties while nodes can belong to multiple communities simultaneously in real-world social as well as other types of networks (for example, a movie can be classified into multiple genres, individuals belong to multiple social groups, scientists may belong to more than one research group, and proteins have multiple functions) palla2005uncovering ; xie2013overlapping ; javed2018community . To address this limitation, the Mixed Membership Stochastic Blockmodel (MMSB) MMSB and the Degree-Corrected Mixed Membership (DCMM) model MixedSCORE extend SBM and DCSBM to overlapping un-weighted networks, where DCMM can also be viewed as an extension of MMSB by considering node heterogeneity. In recent years, some spectral algorithms with theoretical guarantees on consistency have been proposed to study MMSB and DCMM, to name a few, MaoSVM ; mao2020estimating ; MixedSCORE . Thorough survey of the community detection approaches for un-weighted networks has been provided by jin2021survey .
Edge weights are commonplace in complex networks and they provide more information for community detection power than edges newman2004analysis . Edge weights usually tells us how strong or weak a relationship is. For example, in a coauthorship network Newman2004 , edge weight between two authors is the number of papers they co-authored. In a signed network ferligoj1996analysis ; tang2016survey ; kirkley2019balance ; shi2019dynamics , positive edges represent friendship/trust and negative edges mean animosity/mistrust. Three examples of real-world weighted networks in which edge weights indicate the strength of relationship are shown in Figure 1. However, the aforementioned models ignore edge weights and they only model un-weighted networks. To address this limitation, some statistical models are developed in recent years. For non-overlapping weighted networks, several Weighted Stochastic Blockmodels (WSBMs) aicher2015learning ; jog2015information ; peixoto2018nonparametric ; palowitch2018significance ; xu2020optimal ; ng2021weighted , the Distribution-Free model (DFM) qing2021DFM , and the Degree-Corrected Distribution-Free model (DCDFM) DCDFM , and the bipartite versions of DFM and DCDFM BiDFMs extend SBM and DCSBM to weighted networks by allowing edge weights to follow distributions more than Bernoulli distribution. However, the above models for weighted networks assume that each node belongs to a single community and they can not capture the community structure of overlapping weighted networks. Though the weighted version of MMSB (WMMSB) dulac2020mixed and the multi-way blockmodels (MWB) airoldi2013multi can model overlapping weighted networks, they both face some limitations. WMMSB requires edge weights to be generated from Poisson distribution, so it fails to model un-weighted networks, signed networks, networks with positive and negative edge weights, and networks with decimal edge weights. MWB requires edge weights to follow Normal or Bernoulli distributions, thus it fails to describe the community structure of signed networks and networks in which edge weights are nonnegative integers. Meanwhile, both WMMSB and MWB ignore the variation of node degrees. To close these limitations of WMMSB and MWB, we aim at providing a new model for overlapping weighted networks. The main contributions of this article are summarized as follows.
(a) Slovene Parliamentary Party ferligoj1996analysis
(b) Gahuku-Gama subtribes read1954cultures
(c) Karate-club-weighted zachary1977information
Figure 1: Illustrative examples of three real-world weighted networks. For the Slovene Parliamentary Party network, edge weight means political space distance between parties. For the Gahuku-Gama subtribes network, edge weight represents friendship. For the Karate-club-weighted network, edge weight indicates the relative strength of the associations. For visualization, we do not show node labels.
  • 1.
    We provide a generative model called Degree-Corrected Mixed Membership Distribution-Free (DCMMDF) model for overlapping networks with weighted edges. DCMMDF can be seen as an extension of some previous models. In particular, the aforementioned model DCMM is a sub-model of DCMMDF and overlapping signed networks can also be modeled by DCMMDF.
  • 2.
    We use a spectral algorithm with a theoretical guarantee of consistency to fit DCMMDF. The theoretical error rate for the algorithm when edge weights follow a specific distribution can be got immediately from the main result and we also provide several examples to show this.
  • 3.
    We provide overlapping weighted modularity in Equation (7) to measure the quality of overlapping community detection for weighted networks with positive and negative edge weights.
  • 4.
    We provide a method to infer the number of communities for weighted networks by combining the algorithm with the overlapping weighted modularity.
  • 5.
    We implement experiments on computer-generated overlapping weighted networks and eleven real-world networks to demonstrate the advantages of the proposed model and the overlapping weighted modularity.
The rest of the paper is organized as follows. Section 2 introduces the model. Section 3 introduces the algorithm fitting the model. Section 4 presents consistency of the algorithm under the model and provides some examples for different distributions. Section 5 introduces the overlapping weighted modularity and the method to determine the number of communities. Section 6 conducts extensive experiments. Section 7 concludes.
Notations. We use the following notations in this article. Write $[m]:=\{1,2,\ldots,m\}$ for any positive integer $m$. For a vector $x$ and fixed $q>0$, $\|x\|_{q}$ means its $l_{q}$-norm. For a matrix $M$, $M^{\prime}$ means its transpose, $\|M\|_{F}$ means its Frobenius norm, and $\|M\|_{2\rightarrow\infty}$ means the maximum $l_{2}$-norm among all rows of $M$. Let $\mathrm{rank}(M)$ be $M$’s rank. Let $\sigma_{k}(M)$ be $M$’s $k$-th largest singular value, $\lambda_{k}(M)$ be $M$’s $k$-th largest eigenvalue in magnitude, and $\kappa(M)$ be $M$’s condition number. Let $M(i,:)$ (and $M(:,j)$) be the $i$-th row (and the $j$-th column) of matrix $M$. $M(S_{r},:)$ and $M(:,S_{c})$ denote the rows and columns in the index sets $S_{r}$ and $S_{c}$ of matrix $M$, respectively. For any matrix $M$, let $Y=\mathrm{max}(0,M)$ be the matrix such that $Y_{ij}=\mathrm{max}(0,M_{ij})$ for any $i,j$. Let $e_{i}$ be the indicator vector with a $1$ in entry $i$ and 00 in all others.
2 The Degree-Corrected Mixed Membership Distribution-Free model
Given a weighted network $\mathcal{N}$ with $n$ nodes (network is also known as graph and node is also known as a vertex in literature), let $A\in\mathbb{R}^{n\times n}$ be its adjacency matrix such that $A(i,j)$ denotes the edge weight between nodes $i$ and $j$ for $i,j\in[n]$. Since we only consider un-directed networks in this article, $A$ is symmetric. All entries of $A$ are allowed to be any finite real values in this article. We assume there are $K$ communities
$\mathcal{C}_{(1)},\mathcal{C}_{(2)},\ldots,\mathcal{C}_{(K)}.$ (1)
We work with the mixed membership (also known as overlapping) weighted network in which a node can belong to multiple communities. Let $\Pi$ be an $n\times K$ matrix such that $\Pi(i,k)$ denotes the weight of node $i$ belonging to community $k$ for $i\in[n],k\in[K]$. We assume that $\Pi$ should satisfy the following conditions to make our model identifiable:
$\mathrm{rank}(\Pi)=K,\Pi(i,k)\geq 0,\sum_{k=1}^{K}\Pi(i,k)=1\mathrm{~{}for~{}}i\in[n],k\in[K],$ (2)
$\mathrm{Each~{}of~{}the~{}}K\mathrm{~{}communities~{}has~{}at~{}least~{}one~{}pure~{}node},$ (3)
where node $i$ is a pure node if one element of $\Pi(i,:)$ is 1 and it is a mixed node otherwise. Let $\mathcal{I}$ be an index set of nodes corresponding to $K$ pure nodes, one from each community. Similar to mao2020estimating , we let $\Pi(\mathcal{I},:)$ be the $K\times K$ identity matrix for convenience. Let $P$ be a $K\times K$ block matrix (we also call $P$ connectivity matrix occasionally in this article) and satisfy the following conditions:
$P=P^{\prime},P\in\mathbb{R}^{K\times K},~{}\mathrm{rank}(P)=K,\mathrm{and~{}}P\mathrm{~{}has~{}unit~{}diagonals}.$ (4)
Note that since we consider weighted networks and $A$’s entries can by any finite real values, $P$ can have negative elements and $P$ is not a probability matrix like that of traditional models SBM SBM , DCSBM DCSBM , MMSB MMSB , and DCMM MixedSCORE for un-weighted networks. For convenience, let $P_{\mathrm{max}}=\mathrm{max}_{k,l\in[K]}|P(k,l)|$. Note that $P_{\mathrm{max}}$ can be larger than 1 because we will consider node heterogeneity later. Unless specified, throughout this article, $K$ is assumed to be a known integer.
Introduce the $n\times 1$ vector $\theta$ such that $\theta(i)$ is the node heterogeneity of node $i$ for $i\in[n]$, where $\theta$ controls the degree variations for all nodes. Let $\Theta$ be an $n\times n$ diagonal matrix whose $i$-th diagonal element is $\theta(i)$. Set $\theta_{\mathrm{max}}=\mathrm{max}_{i\in[n]}\theta(i)$ and $\theta_{\mathrm{min}}=\mathrm{min}_{i\in[n]}\theta(i)$. Node heterogeneity should be positive, so we have
$\theta_{\mathrm{min}}>0.$ (5)
For arbitrary distribution $\mathcal{F}$, our model generates $A$ for the weighted network $\mathcal{N}$ by assuming that $A(i,j)$ are independent random variables generated from $\mathcal{F}$ with expectation
$\mathbb{E}[A(i,j)]=\Omega(i,j),\mathrm{where~{}}\Omega:=\Theta\Pi P\Pi^{\prime}\Theta\mathrm{~{}for~{}}i,j\in[n].$ (6)
Equation (6) means that $A$’s expectation matrix $\Omega$ should have a block structure $\Theta\Pi P\Pi^{\prime}\Theta$ reflecting community membership information, and this is the reason we call $P$ a block matrix. Meanwhile, Equation (6) also means that $A$ is a random matrix generated from any distribution $\mathcal{F}$ with expectation matrix $\Omega$ and there is no limitation on $\mathcal{F}$ as long as Equation (6) holds.
Definition 1.
Call Equations (1)-(6) the Degree-Corrected Mixed Membership Distribution-Free (DCMMDF) model and denote it by $DCMMDF_{n}(K,P,\Pi,\Theta,\mathcal{F})$.
{tikzpicture}
[…]
{tikzpicture}
[…]
{tikzpicture}
[…]
{tikzpicture}
[…]
Figure 2: Illustrations for overlapping (and non-overlapping) networks can be modeled by DCMMDF. Panel (a): un-weighted network. Panel (b): weighted network with weights being nonnegative integers. Panel (c): weighted network with positive and negative weights. Panel (d): signed network.
Under DCMMDF, node $i$’s expectation degree is $\sum_{j=1}^{n}\Omega(i,j)=\theta(i)\Pi(i,:)\sum_{j=1}^{n}P\Pi^{\prime}(j,:)$. We see that if there is a node $\bar{i}$ such that $\Pi(\bar{i},:)=\Pi(i,:)$ (i.e., node $\bar{i}$ has the same community membership as node $i$) for $\bar{i}\neq i$, the expectation degree of node $\bar{i}$ may still differ from that of node $i$ if $\theta(\bar{i})\neq\theta(i)$, and this is the reason we call $\theta(i)$ heterogeneity parameter. For visualization, Figure 2 sketches some overlapping (and non-overlapping) networks that can be modeled by our DCMMDF.
Remark 1.
By Proposition A.1 of MixedSCORE , DCMMDF is identifiable when Equations (2)-(4) hold, where Proposition A.1 of MixedSCORE has no constraint on $A$’s distribution and it allows $P$ to have negative elements as long as $P$ has unit-diagonals. However, if $P$ has non-unit diagonals, DCMMDF is unidentifiable unless all entries of $\theta$ are the same.
Remark 2.
In this remark, we compare our DCMMDF with some previous models.
  • 1.
    When $\mathcal{F}$ is Bernoulli distribution, DCMMDF reduces to DCMM MixedSCORE , i.e., DCMMDF extends DCMM from un-weighted networks to weighted networks and DCMM is a sub-model of DCMMDF.
  • 2.
    When all nodes are pure, DCMMDF reduces to DCDFM DCDFM , i.e., DCMMDF extends DCDFM from non-overlapping weighted networks to overlapping weighted networks.
  • 3.
    When all nodes are pure and all entries of $\theta$ are the same, DCMMDF reduces to DFM qing2021DFM .
  • 4.
    When $\mathcal{F}$ is Bernoulli distribution and all entries of $\theta$ are the same, DCMMDF reduces to MMSB MMSB .
  • 5.
    When $\mathcal{F}$ is Bernoulli distribution and all nodes are pure, DCMMDF reduces to DCSBM DCSBM .
  • 6.
    When $\mathcal{F}$ is Bernoulli distribution, all nodes are pure, and all entries of $\theta$ are the same, DCMMDF reduces to SBM SBM .
  • 7.
    When $\mathcal{F}$ is Bernoulli distribution, $K$ is 1, and all entries of $\theta$ are the same, DCMMDF reduces to the Erdös-Rényi random graph erdos1960evolution .
After designing our model DCMMDF, a random adjacency matrix $A$ with community membership $\Pi$ can be generated from DCMMDF by the following steps.
  • Step (a)
    Fix $n,K,\Pi,P$, and $\Theta$ satisfying conditions in Equations (2)-(5).
  • Step (b)
    Compute $\Omega=\Theta\Pi P\Pi^{\prime}\Theta$ by Equation (6).
  • Step (c)
    For $1\leq i\leq j\leq n$, let $A(i,j)$ be a random variable generated from distribution $\mathcal{F}$ with expectation $\Omega(i,j)$. Set $A(j,i)=A(i,j)$ since we only consider un-directed weighted networks. If we do not consider self-connected nodes, let $A$’s diagonal entries be zeros.
After generating the random adjacency matrix $A$ by Steps (a)-(c), in this article, we aim at answering the following questions:
  • Q (1)
    Can we design an algorithm to estimate $\Pi$ with known $A$ and $K$ when $A$ is generated from arbitrary distribution $\mathcal{F}$ satisfying Equation (6) under our DCMMDF?
  • Q (2)
    When there exists an algorithm to estimate $\Pi$ such that Q (1) is solved, let $\hat{\Pi}$ be the estimated membership matrix. For any distribution $\mathcal{F}$ satisfying Equation (6), does $\hat{\Pi}$ concentrate around $\Pi$ or can we obtain a theoretical upper bound of the difference between $\hat{\Pi}$ and $\Pi$? Does the upper bound enjoy consistency such that it goes to zero as the network size $n$ goes to infinity?
  • Q (3)
    Can we design a method to determine the number of communities $K$ for weighted networks generated from DCMMDF?
In the next few sections, we will answer the above questions by applying an efficient spectral algorithm to fit DCMMDF, building a theoretical guarantee of consistency for the algorithm under DCMMDF, and proposing overlapping weighted modularity to infer $K$.
3 Algorithm
In this section, we aim at answering Q (1) by designing an efficient spectral algorithm to fit DCMMDF. Because $\Omega$ has a low dimensional structure with rank $K$ when $\mathrm{rank}(\Pi)=K,\mathrm{rank}(P)=K$, and all diagonal entries of $\Theta$ are positive, to answer Q1, one natural way to design an algorithm to fit DCMMDF is from the oracle case with known $\Omega$. A good algorithm fitting DCMMDF should satisfy the following nice property: it returns $\Pi$ exactly when using $\Omega$ to replace $A$. Otherwise, if the algorithm can not return $\Pi$ when the input matrix is $\Omega$, it is hard to say that it fits the model well.
To design an algorithm from the oracle case, because $\Omega$ has a low dimensional structure with $K$ nonzero eigenvalues, we let $\Omega=U\Lambda U^{\prime}$ be the eigendecomposition of $\Omega$ such that $U\in\mathbb{R}^{n\times K},\Lambda\in\mathbb{R}^{K\times K}$, and $U^{\prime}U=I_{K}$, where $\Lambda$’s $k$-th diagonal entry is the leading $k$ eigenvalue of $\Omega$ for $k\in[K]$. Let $U_{*}$ be an $n\times K$ matrix such that its $i$-th row is $U_{*}(i,:)=\frac{U(i,:)}{\|U(i,:)\|_{F}}$ for $i\in[n]$. Let $N_{U}$ be an $n\times n$ diagonal matrix such that its $i$-th diagonal entry is $N_{U}(i,i)=\frac{1}{\|U(i,:)\|_{F}}$ for $i\in[n]$, i.e., we have $U_{*}=N_{U}U$. Lemma 3 of Criterion is a distribution-free theoretical result and it shows that there exists a matrix $Y\in\mathbb{R}^{n\times K}_{+}$ such that $U_{*}=YU_{*}(\mathcal{I},:)$, where $Y=N_{M}\Pi\Theta^{-1}(\mathcal{I},\mathcal{I})N^{-1}_{U}(\mathcal{I},\mathcal{I})$ and $N_{M}$ is a diagonal matrix with all diagonal elements being positive. The structure $U_{*}=YU_{*}(\mathcal{I},:)$ is known as Ideal Cone MaoSVM . Set $Z_{*}=UU^{-1}_{*}(\mathcal{I},:)\sqrt{\mathrm{diag}(U_{*}(\mathcal{I},:)\Lambda U^{\prime}_{*}(\mathcal{I},:))}$. When Equation (4) holds such that $P$ has unit-diagonal elements, Equation (8) of Criterion says that $\Pi(i,:)=\frac{Z_{*}(i,:)}{\|Z_{*}(i,:)\|_{1}}$ for $i\in[n]$, i.e., we can recover $\Pi$ from $Z_{*}$ as long as we know the index set $\mathcal{I}$ when $\Omega$ is given, where $Z_{*}$ is directly related to the eigendecomposition of $\Omega$. Since $U_{*}=YU_{*}(\mathcal{I},:)$ is an Ideal Cone structure, the SVM-cone algorithm designed in MaoSVM can be applied to exactly recover $U_{*}(\mathcal{I},:)$ when the inputs are $U_{*}$ and $K$.
Given $\Omega$ and $K$, we provide the Ideal ScD algorithm below, where S denotes SVM, c denotes cone, and D denotes DCMMDF. Input $\Omega,K$. Output: $\Pi$.
  • 1.
    Compute the leading $K$ eigendecomposition of $\Omega$ to get $U$ and $\Lambda$. Get $U_{*}$ from $U$.
  • 2.
    Apply SVM-cone algorithm MaoSVM to $U_{*}$ with $K$ communities to get the index set $\mathcal{I}$.
  • 3.
    Set $Z_{*}=UU^{-1}_{*}(\mathcal{I},:)\sqrt{\mathrm{diag}(U_{*}(\mathcal{I},:)\Lambda U^{\prime}_{*}(\mathcal{I},:))}$.
  • 4.
    Set $\Pi(i,:)=\frac{Z_{*}(i,:)}{\|Z_{*}(i,:)\|_{1}}$ for $i\in[n]$.
Since the SVM-cone algorithm exactly returns $U_{*}(\mathcal{I},:)$, the Ideal ScD exactly recovers $\Pi$, and this also guarantees the identifiability of our DCMMDF.
For the real case when we only know $A$ and $K$, we aim at estimating $\Pi$ when $A$ is generated from any distribution $\mathcal{F}$ satisfying Equation (6) under our DCMMDF. Algorithm 1, which we call ScD, is a natural extension of the Ideal ScD.
Algorithm 1 ScD
1:The adjacency matrix $A\in\mathbb{R}^{n\times n}$ and the number of communities $K$.
2:The estimated $n\times K$ membership matrix $\hat{\Pi}$.
3:Let $\tilde{A}=\hat{U}\hat{\Lambda}\hat{U}^{\prime}$ be the leading $K$ eigendecomposition of $A$ such that $\hat{\Lambda}$ contains $A$’s leading $K$ eigenvalues. Let $\hat{U}_{*}$ be the row-normalization of $\hat{U}$ such that $\hat{U}_{*}(i,:)=\frac{\hat{U}(i,:)}{\|\hat{U}(i,:)\|_{F}}$ for $i\in[n]$.
4:Apply SVM-cone algorithm MaoSVM to $\hat{U}_{*}$ with $K$ communities to get the estimated index set $\hat{\mathcal{I}}$.
5:Set $\hat{Z}_{*}=\mathrm{max}(0,\hat{U}\hat{U}^{-1}_{*}(\hat{\mathcal{I}},:)\sqrt{\mathrm{diag}(\hat{U}_{*}(\hat{\mathcal{I}},:)\hat{\Lambda}\hat{U}^{\prime}_{*}(\hat{\mathcal{I}},:))})$.
6:Estimate $\Pi(i,:)$ by setting $\hat{\Pi}_{*}(i,:)=\hat{Z}_{*}(i,:)/\|\hat{Z}_{*}(i,:)\|_{1}$ for $i\in[n]$.
Remark 3.
To obtain binary rather fuzzy community membership, we can threshold each entry of $\hat{\Pi}$ to obtain a binary membership matrix $\hat{\Pi}^{0}\in\{0,1\}^{n\times K}$ such that
[…]
where $\delta_{K}$ is a threshold and we set it as $\frac{1}{K}$ in this paper. Follow a similar definition of $\hat{\Pi}^{0}$, we can define $\Pi^{0}\in\{0,1\}^{n\times K}$ from $\Pi$. The two binary membership matrices will be used in defining our overlapping weighted modularity.
Because $A$’s expectation is $\Omega$ under the DCMMDF, $\hat{U},\hat{\Lambda},\hat{Z}_{*}$, and $\hat{\Pi}$ are good estimations of $U,\Lambda,Z_{*}$, and $\Pi$, respectively. Again, ScD is the SVM-cone-DCMMSB algorithm of MaoSVM and we re-name the SVM-cone-DCMMSB as ScD to emphasize its distribution-free property. Note that in MaoSVM , the ScD algorithm is only used to estimate $\Pi$ when $A$ is generated from a Bernoulli distribution under DCMM while the ScD used in this article can estimate $\Pi$ when $A$ is generated from any distribution $\mathcal{F}$ as long as Equation (6) holds under our DCMMDF. Furthermore, building spectral algorithms via applications of the eigendecomposition of the adjacency matrix or the Laplacian matrix to fit statistical models for un-weighted networks is a popular way in community detection areas, see rohe2011spectral ; lei2015consistency ; joseph2016impact ; SCORE ; MixedSCORE ; MaoSVM ; mao2020estimating .
The computationally intensive steps of ScD are the leading $K$ eigendecomposition (i.e., SVD) and SVM-cone. The computational cost of SVD is $O(n^{3})$. Because $A$ for real-world networks is usually sparse, applying the power method newman2006modularity , the computation complexity for SVD is $O(Kn^{2})$ SCORE ; MixedSCORE . The complexity of the SVM-cone algorithm is $O(nK^{2})$ chang2011libsvm ; MaoSVM . Since $K\ll n$ in this article, as a result, the total complexity of ScD is $O(Kn^{2})$.
4 Asymptotic Consistency
In this section, we aim at answering Q (2) by building a theoretical guarantee on ScD’s consistency for any distribution $\mathcal{F}$ satisfying Equation (6) under the DCMMDF model. Set $\tau=\mathrm{max}_{i,j\in[n]}|A(i,j)-\Omega(i,j)|$ and $\gamma=\mathrm{max}_{i,j\in[n]}\frac{\mathrm{Var}(A(i,j))}{\theta(i)\theta(j)}$, where $\mathrm{Var}(A(i,j))$ denotes the variance of $A(i,j)$. $\tau$ and $\gamma$ are two parameters related to distribution $\mathcal{F}$. For different distributions $\mathcal{F}$, the upper bounds of $\tau$ and $\gamma$ can be different, see Examples 1-5 for detail. We need the following assumption for our theoretical analysis.
Assumption 1.
Under $DCMMDF_{n}(K,P,\Pi,\Theta,\mathcal{F})$, assume $\gamma\theta_{\mathrm{max}}\|\theta\|_{1}\geq\tau^{2}\mathrm{log}(n)$.
Assumption 1 provides a lower bound requirement on $\gamma\theta_{\mathrm{max}}\|\theta\|_{1}$ for our theoretical analysis and it functions similar to Assumption 3.1 of mao2020estimating and Assumption 1 of MixedSCORE . We also need the following condition which simplifies our theoretical result.
Condition 1.
Under $DCMMDF_{n}(K,P,\Pi,\Theta,\mathcal{F})$, assume $\kappa(P)=O(1),K=O(1),\lambda_{K}(\Pi^{\prime}\Pi)=O(\frac{n}{K}),\theta_{\mathrm{max}}=O(\sqrt{\rho})$, and $\theta_{\mathrm{min}}=O(\sqrt{\rho})$ for some $\rho>0$.
In Condition 1, $\kappa(P)=O(1)$ is mild because it means the connectivity matrix $P$ is well-conditioned, $K=O(1)$ means we only consider a constant number of communities for our theoretical analysis, $\lambda_{K}(\Pi^{\prime}\Pi)=O(\frac{n}{K})$ means the size of each community is in the same order, and $\theta_{\mathrm{max}}=O(\sqrt{\rho}),\theta_{\mathrm{min}}=O(\sqrt{\rho})$ mean that all entries of $\theta$ are in the same order. So, we see that Condition 1 is mild. We are ready to present the main theorem:
Theorem 1.
Under $DCMMDF_{n}(K,P,\Pi,\Theta,\mathcal{F})$, let $\hat{\Pi}$ be obtained from Algorithm 1, when Assumption 1 and Condition 1 hold, there exists a permutation matrix $\mathcal{P}\in\mathbb{R}^{K\times K}$ such that with probability at least $1-o(n^{-5})$, we have
$\mathrm{max}_{i\in[n]}\|e^{\prime}_{i}(\hat{\Pi}-\Pi\mathcal{P})\|_{1}=O(\frac{\sqrt{\gamma\mathrm{log}(n)}}{\sigma_{K}(P)\sqrt{\rho n}}).$
Proof of Theorem 1 is provided in A. In Theorem 1, when $P,K,\theta_{\mathrm{min}}$, and distribution $\mathcal{F}$ are fixed, increasing $n$ decreases ScD’s error rate. In particular, when $n$ goes to infinity, the theoretical upper bound of the error rate in Theorem 1 goes to zero, and this guarantees ScD’s consistency under our DCMMDF. In Theorem 1, only parameters $\gamma$ and $\tau$ are directly related to distribution $\mathcal{F}$, so our main theoretical result in Theorem 1 is general since we do not constrain $A$’s distribution $\mathcal{F}$. In the following examples, we will show that $\gamma$ is finite and we can obtain a more accurate theoretical upper bound of ScD’s error rate by analyzing $\gamma$’s upper bound for some specific distributions.
Example 1.
When $\mathcal{F}$ is Normal distribution such that $A(i,j)\sim\mathrm{Normal}(\Omega(i,j),\sigma^{2}_{A})$, i.e., $A(i,j)\in\mathbb{R}$, we have $\mathbb{E}[A(i,j)]=\Omega(i,j)$ satisfying Equation (6), $P$’s non-diagonal elements can be any finite real values because the mean of Normal distribution can be negative, $\Omega(i,j)$’s range is (,)($\infty$,$\infty$)( - ∞ , ∞ ), $\tau$ is an unknown finite positive real value, and $\gamma=\mathrm{max}_{i,j\in[n]}\frac{\mathrm{Var}(A(i,j))}{\theta(i)\theta(j)}\leq\frac{\sigma^{2}_{A}}{\theta^{2}_{\mathrm{min}}}=O(\frac{\sigma^{2}_{A}}{\rho})$ is finite. Setting $\gamma=O(\frac{\sigma^{2}_{A}}{\rho})$ in Theorem 1 obtains the theoretical upper bound of ScD’s error rate when $\mathcal{F}$ is Normal distribution. We see that increasing $\rho$ (or decreasing $\sigma^{2}_{A}$) decreases ScD’s error rate for Normal distribution.
Example 2.
When $\mathcal{F}$ is Bernoulli distribution such that $A(i,j)\sim\mathrm{Bernoulli}(\Omega(i,j))$ and our DCMMDF degenerates to the DCMM model, i.e., $A(i,j)\in\{0,1\}$, we have $\mathbb{E}[A(i,j)]=\Omega(i,j)$ satisfying Equation (6), all entries of $P$ should be nonnegative, $\Omega(i,j)$ is a probability in $[0,1]$, and $\tau=1$. For this case, we have $\mathrm{Var}(A(i,j))=\Omega(i,j)(1-\Omega(i,j))\leq\Omega(i,j)=\theta(i)\theta(j)\Pi(i,:)P\Pi^{\prime}(j,:)\leq\theta(i)\theta(j)P_{\mathrm{max}}$, which suggests that $\gamma=P_{\mathrm{max}}$, a finite number. Setting $\gamma=P_{\mathrm{max}}$ in Theorem 1 gets the theoretical upper bound of ScD’s error rate and we find that increasing $\rho$ decreases ScD’s error rate. Setting $\gamma=P_{\mathrm{max}}$ and $\tau=1$, Assumption 1 is $\rho n\geq O(\mathrm{log}(n))$ by Condition 1, and it means a lower bound requirement on network sparsity. So $\rho$ controls network sparsity for Bernoulli distribution.
Example 3.
When $\mathcal{F}$ is Poisson distribution such that $A(i,j)\sim\mathrm{Poisson}(\Omega(i,j))$, i.e., $A(i,j)\in\mathbb{N}$, where $\mathbb{N}$ is the set of nonnegative integers, we have $\mathbb{E}[A(i,j)]=\Omega(i,j)$ satisfying Equation (6), $P$’s elements should be nonnegative, $\Omega(i,j)$’s range is $[0,\infty)$, $\tau$ is an unknown positive integer, and $\gamma=\mathrm{max}_{i,j\in[n]}\frac{\mathrm{Var}(A(i,j))}{\theta(i)\theta(j)}=\mathrm{max}_{i,j\in[n]}\frac{\Omega(i,j)}{\theta(i)\theta(j)}\leq P_{\mathrm{max}}$, i.e., $\gamma$ is finite. Setting $\gamma=P_{\mathrm{max}}$ in Theorem 1 gets the theoretical result for Poisson distribution and we find that increasing $\rho$ decreases ScD’s error rate.
Example 4.
When $\mathcal{F}$ is Uniform distribution such that $A(i,j)\sim\mathrm{Uniform}(0,2\Omega(i,j))$ for $i,j\in[n]$, i.e., $A(i,j)\in(0,2\rho)$, we have $\mathbb{E}[A(i,j)]=\frac{0+2\Omega(i,j)}{2}=\Omega(i,j)$ satisfying Equation (6), $P$’s entries should be nonnegative, $\Omega(i,j)$’s range is $[0,+\infty)$, $\tau$ is an unknown value with a upper bound $2\rho$, and $\gamma=\mathrm{max}_{i,j\in[n]}\frac{\mathbb{E}[(A(i,j)-\Omega(i,j))^{2}]}{\theta(i)\theta(j)}=\mathrm{max}_{i,j\in[n]}\frac{4\Omega^{2}(i,j)}{12\theta(i)\theta(j)}\leq\mathrm{max}_{i,j\in[n]}\frac{(\theta(i)\theta(j)P_{\mathrm{max}})^{2}}{3\theta(i)\theta(j)}=O(\frac{\rho P^{2}_{\mathrm{max}}}{3})$. Setting $\gamma=O(\frac{\rho P^{2}_{\mathrm{max}}}{3})$ in Theorem 1, we find that $\rho$ disappears in the theoretical upper bound of ScD’s error rate, which indicates that increasing $\rho$ has almost no significant impact on the behavior of ScD.
Example 5.
DCMMDF can also model overlapping signed networks by setting $\mathbb{P}(A(i,j)=1)=\frac{1+\Omega(i,j)}{2}$ and $\mathbb{P}(A(i,j)=-1)=\frac{1-\Omega(i,j)}{2}$, i.e., $A(i,j)\in\{1,-1\}$ for $i\neq j$. For signed networks, $\mathbb{E}[A(i,j)]=\Omega(i,j)$ holds by basic computation, $P$’s non-diagonal elements can be negative, $\Omega(i,j)$’s range is $[-1,1]$ because $\frac{1+\Omega(i,j)}{2}$ and $\frac{1-\Omega(i,j)}{2}$ are probabilities, $\tau$ is $2$, and $\mathrm{Var}(A(i,j))=1-\Omega^{2}(i,j)\leq 1$, i.e., $\gamma\leq\frac{1}{\theta^{2}_{\mathrm{min}}}=O(\frac{1}{\rho})$ is finite. Setting $\gamma=O(\frac{1}{\rho})$ in Theorem 1 obtains theoretical results for the signed network and we see that increasing $\rho$ decreases ScD’s error rate.
Generally speaking, benefiting from the distribution-free property of our DCMMDF, $\mathcal{F}$ can be any distribution as long as Equation (6) holds such as the Binomial, Double exponential, Gamma, Exponential, and Laplace distributions in http://www.stat.rice.edu/~dobelman/courses/texts/distributions.c&b.pdf.
By Examples 1, 4, and 5, we find that the non-diagonal elements of $A$ are always nonzero, which suggests that there is always a link between any two nodes. However, this is impractical for real-world large-scale networks in which most node pairs have no connections lei2015consistency . Treat an edge with weight 0 as a missing edge. Following the idea in qing2021DFM , we can generate missing edges for networks generated from DCMMDF such that $A(i,j)$ is multiplied by $\mathcal{A}(i,j)$ for $i,j\in[n]$, where $\mathcal{A}\in\{0,1\}^{n\times n}$ is a symmetric and connected adjacency matrix for an un-weighted network and it can be generated from any model for un-directed and un-weighted networks such as SBM, DCSBM, MMSB, DCMM, and Erdös-Rényi random graph.
5 Estimation of the Number of Communities
In this section, we aim at answering Q (3). Algorithm 1 requires a known number of communities $K$. However, for real-world networks, $K$ is usually unknown, which diminishes the usefulness of a community detection approach. Here, to estimate $K$, we introduce overlapping weighted modularity.
Because $A$ considered in this article may have negative elements, $A$ can be written as $A=A^{+}-A^{-}$, where $A^{+}=\mathrm{max}(0,A)$ and $A^{-}=\mathrm{max}(0,-A)$. Let $d^{+}$ be an $n\times 1$ vector such that $d^{+}(i)=\sum_{j=1}^{n}A^{+}(i,j)$ for $i\in[n]$. Let $m^{+}=\sum_{i=1}^{n}d^{+}_{i}$. Similarly, we can define $d^{-}$ and $m^{-}$ from $A^{-}$. Let $O$ be an $n\times 1$ vector such that $O(i)$ denotes the number of communities that node $i$ belongs to, i.e., $O(i)=\sum_{k\in[K]}\hat{\Pi}^{0}(i,k)$ for $i\in[n]$, where $\hat{\Pi}^{0}$ is a binary membership matrix computed by Remark 3. Let $\hat{\mathcal{C}}_{k}=\{i:\hat{\Pi}^{0}(i,k)=1,i\in[n]\}$ be the set of nodes belonging to community $k$ for $k\in[K]$. For convenience, let $\hat{\mathcal{C}}$ be the set of the $K$ estimated communities $\hat{\mathcal{C}}_{k}$ for $k\in[K]$. Based on the binary membership matrix $\hat{\Pi}^{0}$, overlapping modularity $Q^{+}$ for $A$’s positive entries and overlapping modularity $Q^{-}$ for $A$’s negative entries are defined as
$Q^{+}=\frac{1_{m^{+}>0}}{m^{+}}\sum_{k=1}^{K}\sum_{i\in\hat{\mathcal{C}}_{k},j\in\hat{\mathcal{C}}_{k}}\frac{1}{O(i)O(j)}(A^{+}(i,j)-\frac{d^{+}(i)d^{+}(j)}{m^{+}}),Q^{-}=\frac{1_{m^{-}>0}}{m^{-}}\sum_{k=1}^{K}\sum_{i\in\hat{\mathcal{C}}_{k},j\in\hat{\mathcal{C}}_{k}}\frac{1}{O(i)O(j)}(A^{-}(i,j)-\frac{d^{-}(i)d^{-}(j)}{m^{-}}),$
where $1_{m^{+}>0}$ is indicator function such that $11_{m^{+}>0}=$ 1 if $m^{+}>0$ and 0 otherwise. Similar for $1_{m^{-}>0}$. Following the idea of defining weighted modularity for signed networks gomez2009analysis , our overlapping weighted modularity is defined as
$Q=\frac{m^{+}}{m^{+}+m^{-}}Q^{+}-\frac{m^{-}}{m^{+}+m^{-}}Q^{-}.$ (7)
Our overlapping weighted modularity evaluates the quality of overlapping community partition. Similar to the Newman-Girvan modularity newman2004finding ; newman2006modularity and weighted modularity gomez2009analysis , a larger overlapping weighted modularity indicates a better community partition. Our overlapping weighted modularity includes several popular modularity as special cases.
  • 1.
    When $A$ has both positive and negative entries and $\hat{\Pi}^{0}$ degenerates such that each node only belongs to one community, our overlapping weighted modularity reduces to the modularity for signed networks provided in Equation (17) of gomez2009analysis .
  • 2.
    When all entries of $A$ are nonnegative (i.e., $m^{-}=0$), our overlapping weighted modularity degenerates to the overlapping modularity provided in Equation (2) of shen2009detect .
  • 3.
    When $\hat{\Pi}^{0}$ degenerates such that each node only belongs to one community and all entries of $A$ are nonnegative, our overlapping weighted modularity degenerates to the famous Newman-Girvan modularity newman2004finding ; newman2006modularity .
To infer the number of communities, we adopt the suggestion proposed in nepusz2008fuzzy , that is, we iteratively increase $k$ and choose the one maximizing our overlapping weighted modularity in Equation (7). We let $\hat{\Pi}^{0}$ be obtained from our ScD algorithm and call our method for inferring $K$ as KScD.
6 Experiments
We conduct extensive experiments to validate our theoretical results, demonstrate the effectiveness of ScD in estimating membership matrix $\Pi$, and demonstrate the effectiveness of our overlapping weighted modularity by showing that KScD can infer $K$ for networks generated from DCMMDF.
6.1 Baseline Methods
For the task of estimating memberships, we select three mixed membership community detection approaches as baseline methods of our DCMMDF model.
  • 1.
    GeoNMF GeoNMF adapts nonnegative matrix factorization to estimate mixed memberships for networks generated from MMSB.
  • 2.
    SPACL mao2020estimating is a spectral algorithm for estimating mixed memberships for networks generated from MMSB.
  • 3.
    Mixed-SCORE MixedSCORE is a spectral algorithm for estimating mixed memberships for networks generated from DCMM.
Remark 4.
GeoNMF and Mixed-SCORE may fail to output for the case that some elements in the adjacency matrix $A$ are negative. To make GeoNMF and Mixed-SCORE work when $A$ has negative elements, we make all entries of $A$ nonnegative by adding a sufficiently large constant for these two methods.
For the task of inferring $K$, we compare KScD with the following four methods.
  • 1.
    KSPACL is the method to determine $K$ by finding the number of clusters that maximizes our overlapping weighted modularity when $\hat{\Pi}^{0}$ is obtained from SPACL.
  • 2.
    KMixed-SCORE is the method to infer $K$ by maximizing our overlapping weighted modularity when $\hat{\Pi}^{0}$ is obtained from Mixed-SCORE.
  • 3.
    NB and BHac le2022estimating are two methods for inferring the number of communities and they determine $K$ via non-backtracking matrix and Bethe Hessian matrix, respectively.
Remark 5.
For determining $K$, we do not consider KGeoNMF which is defined similarly to KScD, KSPACL, and KMixed-SCORE because KGeoNMF fails to output for large $k$.
6.2 Evaluation metrics
For the task of estimating memberships, we consider the following metrics depending on the fact that whether the ground truth membership matrix $\Pi$ is known.
Metrics with ground truth. For this case, we introduce three widely used evaluation metrics: Hamming error MixedSCORE and Relative error mao2020estimating .
  • 1.
    Hamming error measures the $l_{1}$ difference between $\Pi$ and $\hat{\Pi}$:
    $\mathrm{Hamming~{}error=}\mathrm{min}_{\mathcal{P}\in\{K\times K\mathrm{~{}permutation~{}matrix}\}}\frac{1}{n}\|\hat{\Pi}-\Pi\mathcal{P}\|_{1}.$
    Hamming error is nonnegative with an upper bound 1, and it is the smaller the better. Note that Theorem 1 provides a theoretical upper bound of ScD’s Hamming error for each node.
  • 2.
    Relative error evaluates the relative $l_{2}$ distance between $\Pi$ and $\hat{\Pi}$:
    $\mathrm{Relative~{}error}=\mathrm{min}_{\mathcal{P}\in\{K\times K\mathrm{~{}permutation~{}matrix}\}}\frac{\|\hat{\Pi}-\Pi\mathcal{P}\|_{F}}{\|\Pi\|_{F}}.$
    A smaller Relative error indicates that $\hat{\Pi}$ is a better estimation of membership matrix $\Pi$.
Metrics without ground truth. For this case, our overlapping weighted modularity in Equation (7) is capable to evaluate the quality of detected communities, and a higher value of overlapping weighted modularity indicates more accurately detected communities.
Accuracy rate. For the task of estimating $K$, we use the Accuracy rate to evaluate the precision of KScD and its competitors, where Accuracy rate is defined as the proportion of times an approach exactly determines $K$ le2022estimating .
6.3 Simulations
For computer-generated networks, we set $n=400$ and $K=3$. Let each community has $n_{0}$ number of pure nodes. For pure node $i$, let $\Pi(i,k)=1$ if $1+(k-1)n_{0}\leq i\leq kn_{0}$ for $i\in[Kn_{0}],k\in[K]$. Set $n_{0}=40$. Let all mixed nodes have four different memberships $(0.1,0.1,0.8),(0.1,0.8,0.1),(0.8,0.1,0.1)$, and $(1/3,1/3,1/3)$, each with $\frac{n-3n_{0}}{4}$ number of nodes. Let $P$’s diagonal elements be 1 and non-diagonal elements be $p$. For Normal distribution and signed networks, $p$ is set to be negative. For other distributions, $p$ is set to be positive. Let $\theta(i)=\rho\times(\mathrm{rand}(1)/2+0.5)$, where $\mathrm{rand}(1)$ is a random value ranged in $(0,1)$. In this way, $n,K,P,\Pi$, and $\Theta$ have been set. Then following Steps (a)-(c), we can generate an adjacency matrix $A$ with true membership matrix $\Pi$ for any distribution $\mathcal{F}$ satisfying Equation (6) under our DCMMDF model. For the task of community detection, applying ScD (and its competitors) to $A$ with $K$ communities obtains the estimated membership matrix $\hat{\Pi}$, and then we can obtain Hamming error and Relative error. For each parameter setting considered in this section, we report the averaged Hamming error (and Relative error) over 50 repetitions for each method. For the task of determining the number of communities, applying KScD (and its competitors) to $A$ obtains the estimated number of communities, and then we can obtain the Accuracy rate over the 50 repetitions.
Remark 6.
For our synthetic networks, $P,\Pi$, and $\Theta$ can be set arbitrarily as long as Equations (2)-(5) hold and $P$’s non-diagonal elements should be positive or nonnegative or can be negative depending on distribution $\mathcal{F}$ just as analyzed in Examples 1-5. Meanwhile, for simulations, $\mathcal{F}$ can be any distribution since our DCMMDF is a distribution-free model and it has no constraint on a specific distribution $\mathcal{F}$ as long as Equation (6) holds. Furthermore, to save space, we do not consider missing edges in the simulations.
Experiment 1: Normal distribution. Set $\mathcal{F}$ as Normal distribution such that $A(i,j)\sim\mathrm{Normal}(\Omega(i,j),\sigma^{2}_{A})$. Let $p=-0.2,\sigma^{2}_{A}=2$, and $\rho$ range in $\{0.1,0.2,\ldots,2\}$. The results are in panels (a)-(c) of Figure 3. We observe the following: (i) ScD performs better as $\rho$ increases, and this is consistent with our theoretical analysis in Example 1. (ii) ScD, GeoNMF, SPACL, and Mixed-SCORE perform similarly and they all successfully estimate memberships for large $\rho$. (iii) For the task of inferring the number of communities, our methods KScD, KSPACL, and KMixed-SCORE designed based on our overlapping weighted modularity enjoy similar and satisfactory performances because their accuracy rates increase as $\rho$ increases, which indicates the effectiveness of our overlapping weighted modularity. For comparison, NB and BHac fail to estimate $K$ for Normal distribution.
Remark 7.
For the task of estimating memberships, in our experimental studies, we aim at using simulations to verify our theoretical analysis that ScD has different behaviors for different distributions when $\rho$ is changed as analyzed in Examples 1-5. We also want to investigate the performances of GeoNMF, SPACL, and Mixed-SCORE when they are applied to estimate memberships for weighted networks generated from our model DCMMDF instead of showing that ScD outperforms its competitors.
Experiment 2: Bernoulli distribution. Set $\mathcal{F}$ as Bernoulli distribution such that $A(i,j)\sim\mathrm{Bernoulli}(\Omega(i,j))$. Let $p=0.2$ and $\rho$ range in $\{0.1,0.1,\ldots,1\}$. Panels (d)-(f) of Figure 3 show the results. For the task of estimating memberships, all methods enjoy better performances as $\rho$ increases which is consistent with our analysis in Example 2. For the task of determining $K$, all methods have larger accuracy rates when $\rho$ increases.
Experiment 3: Poisson distribution. Set $\mathcal{F}$ as Poisson distribution such that $A(i,j)\sim\mathrm{Poisson}(\Omega(i,j))$. Let $p=0.2$ and $\rho$ range in $\{0.2,0.4,\ldots,2\}$. The results displayed in panels (g)-(i) of Figure 3 are similar to that of Experiment 2 and we omit the analysis here.
Experiment 4: Uniform distribution. Set $\mathcal{F}$ as Uniform distribution such that $A(i,j)\sim\mathrm{Uniform}(0,2\Omega(i,j))$. Let $p=0.2$ and $\rho$ range in $\{2,4,\ldots,20\}$. Panels (j)-(l) of Figure 3 display the results. We observe the following: (i) ScD’s performance is almost unchanged when $\rho$ becomes larger, which is consistent with our analysis in Example 4. (ii) ScD, GeoNMF, SPACL, and Mixed-SCORE perform satisfactorily and similarly because their Hamming errors are lesser than 0.07, their Relative errors are lesser than 0.18, and their ONMs are larger than 0.65. (iii) Our KScD, KSPACL, and KMixed-SCORE correctly determine $K$ for this experiment. It is interesting to see that NB and BHac successfully estimate $K$ when $\rho<10$ while they fail to infer $K$ when $\rho\geq 10$.
Experiment 5: Signed network. For signed network when $\mathbb{P}(A(i,j)=1)=\frac{1+\Omega(i,j)}{2}$ and $\mathbb{P}(A(i,j)=-1)=\frac{1-\Omega(i,j)}{2}$. Let $p=-0.2$ and $\rho$ range in $\{0.1,0.2,\ldots,1\}$. The last three panels of Figure 3 display the results. We see that ScD, GeoNMF, SPACL, and Mixed-SCORE perform similarly and they behave better in estimating memberships when $\rho$ increases, which supports our analysis in Example 5. For determining $K$, the accuracy rates of our methods KScD, KSPACL, and KMixed-SCORE increase when increasing $\rho$ while their two competitors fail to find $K$ for signed networks.
(a) Normal distribution
(b) Normal distribution
(c) Normal distribution
(d) Bernoulli distribution
(e) Bernoulli distribution
(f) Bernoulli distribution
(g) Poisson distribution
(h) Poisson distribution
(i) Poisson distribution
(j) Uniform distribution
(k) Uniform distribution
(l) Uniform distribution
(m) Signed networks
(n) Signed networks
(o) Signed networks
Figure 3: Numerical results of Experiments 1-5.
Generally speaking, from the numerical results of Experiments 1-5, we have the following conclusions:
  • 1.
    For the task of estimating memberships, the results support our analysis in Examples 1-5 that ScD performs differently for different distributions when we change $\rho$. GeoNMF, SPACL, and Mixed-SCORE have competitive performances with ScD and they can estimate memberships for networks generated from DCMMDF.
  • 2.
    For the task of determining the number of communities, our methods enjoy satisfactory performances, which suggests the effectiveness of our overlapping weighted modularity. For comparison, NB and BHac fail to estimate $K$ when $\mathcal{F}$ is Normal distribution, Uniform distribution, and signed network.
6.4 Real Data
Table 1: Basic information and summarized statistics of real-world networks studied in this paper.
Dataset Source Node meaning Edge meaning Weighted? True memberships $n$ $K$
Karate club zachary1977information Member Tie No Known 34 2
Dolphins dolphins0 Dolphin Association No Known 62 2
UKfaculty nepusz2008fuzzy Faculty Friendship No Known 79 3
Political books Krebs (unpublished) Book Co-purchase No Known 92 2
Political blogs Polblogs1 Blog Hyperlink No Known 1222 2
Slovene Parliamentary Party ferligoj1996analysis Party Political space distance Yes Unknown 10 2
Gahuku-Gama subtribes read1954cultures Tribe Friendship Yes Known 16 3
Karate-club-weighted zachary1977information Member Tie Yes Known 34 2
Train bombing hayes2006connecting Terrorist Contact Yes Unknown 64 Unknown
Les Misérables knuth1993stanford Character Co-occurence Yes Unknown 77 Unknown
US Top-500 Airport Network colizza2007reaction Airport #Seats Yes Unknown 500 Unknown
In addition to the computer-generated networks, we also consider some real-world datasets. Table 1 displays basic information for real-world networks considered in this article. For the five un-weighted networks in Table 1, their adjacency matrices and true memberships can be downloaded from http://zke.fas.harvard.edu/software/SCOREplus/Matlab/datasets/. For Slovene Parliamentary Party, it can be downloaded from http://vlado.fmf.uni-lj.si/pub/networks/data/soc/Samo/Stranke94.htm. For Gahuku-Gama subtribes, it can be downloaded from http://konect.cc/networks/ucidata-gama/ and its community information is shown in Figure 9 (b) yang2007community . Karate-club-weighted is the weighted version of Karate club and can be downloaded from http://vlado.fmf.uni-lj.si/pub/networks/data/ucinet/ucidata.htm#kazalo. For Train bombing and Les Misérables, they can be downloaded from http://konect.cc/networks/ (see also kunegis2013konect ). For US Top-500 Airport Network, it can be downloaded from https://toreopsahl.com/datasets/#online_social_network.
For the task of determining the number of communities, we apply KScD, KSPACL, KMixed-SCORE, and their competitors to real-world networks in Table 1. Here, for comparison, we also consider another three approaches BHa, BHm, and BHmc designed based on the Bethe Hessian matrix in le2022estimating . Figure 4 displays the overlapping weighted modularity via Equation (7) using ScD (and SPACL and Mixed-SCORE) for different choices of the number of communities. From Figure 4, we can directly find the $k$ maximizing $Q$ for each method. Table 2 shows the estimated number of communities of KScD, KSPACL, KMixed-SCORE, and their competitors for real networks considered in this paper. For networks with known $K$, KScD, KSPCAL, and KMixed-SCORE correctly infer $K$ for all networks except for the Dophins network while NB, BHac, BHa, BHm, and BHmc fail to determine $K$ for Political blogs and all weighted networks. Again, the fact that our methods significantly outperform NB, BHac, BHa, BHm, and BHmc for the task of inferring $K$ on both un-weighted and weighted real-world networks, suggests the effectiveness and advantage of our overlapping weighted modularity in measuring the quality of overlapping community detection. For Train bombing, Les Misérables, and US Top-500 Airport Network, our methods designed based on overlapping weighted modularity infer their $K$ as 2 while $K$ estimated by NB, BHac, BHa, BHm, and BHmc is larger. As a result, for networks with unknown $K$, we tend to believe that $K$ estimated by our KScD, KSPACL, and KMixed-SCORE is more accurate than their competitors because our approaches outperform their competitors for both simulated weighted networks and real-world networks with known $K$.
(a) Karate club
(b) Dolphins
(c) UKfaculty
(d) Political books
(e) Political blogs
(f) Slovene Parliamentary Party
(g) Gahuku-Gama subtribes
(h) Karate-club-weighted
(i) Train bombing
(j) Les Misérables
(k) US Top-500 Airport Network
Figure 4: Overlapping weighted modularity $Q$ computed by Equation (7) against the number of clusters for real data used in this paper.
For the task of estimating memberships, Table 3 reports the overlapping weighted modularity $Q$ of ScD, GeoNMF, SPACL, and Mixed-SCORE. The results show that all methods have similar or the same $Q$ for all networks. In particular, from Table 3, we find that $Q$ of Karate-club-weighted is larger than that of Karate club, which supports the statement in newman2004analysis that weights provide more useful information about community structure than edges.
Table 2: Estimated number of communities by KScD, KSPACL, KMixed-SCORE, and their competitors for real-world networks in Table 1.
Dataset True $K$ KScD KSPACL KMixed-SCORE NB BHac BHa BHm BHmc
Karate club 2 2 2 2 2 2 2 2 2
Dolphins 2 3 3 4 2 2 2 2 2
UKfaculty 3 3 3 3 3 3 3 3 3
Political books 2 2 2 2 2 3 3 2 3
Political blogs 2 2 2 2 7 8 7 7 8
Slovene Parliamentary Party 2 2 2 2 N/A N/A N/A N/A N/A
Gahuku-Gama subtribes 3 3 3 3 1 13 12 1 N/A
Karate-club-weighted 2 2 2 2 4 4 4 4 4
Train bombing Unknown 2 2 2 3 4 4 3 3
Les Misérables Unknown 2 2 2 6 7 7 6 7
US Top-500 Airport Network Unknown 2 2 2 147 158 158 147 147
Table 3: Overlapping weighted modularity $Q$ (calculated by Equation (7) of ScD and its competitors for real-world networks in Table 1. For networks with known $K$, all four methods use the true $K$ to compute $Q$. For networks with unknown $K$, to compute $Q$, all methods use $K$ estimated by KScD in Table 2. For each data, the largest $Q$ is marked in bold.
Dataset ScD GeoNMF SPACL Mixed-SCORE
Karate club 0.3715 0.3715 0.3715 0.3715
Dolphins 0.3848 0.3848 0.3848 0.3899
UKfaculty 0.4384 0.4362 0.4333 0.4393
Political books 0.4786 0.4786 0.4786 0.4786
Political blogs 0.4251 0.4249 0.4249 0.4249
Slovene Parliamentary Party 0.4839 0.4839 0.4839 0.4839
Gahuku-Gama subtribes 0.3825 0.4149 0.4149 0.4310
Karate-club-weighted 0.4036 0.4036 0.4036 0.4036
Train bombing 0.3450 0.2792 0.3450 0.3411
Les Misérables 0.3814 0.3814 0.3814 0.3814
US Top-500 Airport Network 0.2310 0.2297 0.2297 0.2309
More than our overlapping weighted modularity, to have a better understanding of community structure for real data, we define three indices $\zeta_{\mathrm{mixed}},\zeta_{\mathrm{pure}}$, and $\varsigma$ to measure the mixedness, the purity, and the equilibrium of a real network. Call node $i$ a highly mixed node if $\mathrm{max}_{k\in[K]}\hat{\Pi}(i,k)\leq 0.7$ and a highly pure node if $\mathrm{max}_{k\in[K]}\hat{\Pi}(i,k)\geq 0.9$ for $i\in[n]$. $\zeta_{\mathrm{mixed}}$ and $\zeta_{\mathrm{pure}}$ are the proportions of highly mixed nodes and highly pure nodes in a network, respectively. Larger $\zeta_{\mathrm{mixed}}$ (and $\zeta_{\mathrm{pure}}$) means a higher proportion of highly mixed (and pure) nodes in a network. $\varsigma$ is defined as $\varsigma=\frac{\mathrm{min}_{k\in[K]}\|\hat{\Pi}(:,k)\|_{1}}{\mathrm{max}_{k\in[K]}\|\hat{\Pi}(:,k)\|_{1}}$. $\varsigma$ ranges in $(0,1]$ and a larger $\varsigma$ indicates a more balanced network. Table 4 shows the three indices detected by ScD for real data considered in this paper and we have the following conclusions:
  • 1.
    For Karate club, it has $34\times 0.882\approx 3$ highly mixed nodes and $34\times 0.7353\approx 25$ highly pure nodes. This data is quite balanced for its large equilibrium index $\varsigma$.
  • 2.
    For Dolphins, it has $62\times 0.0323\approx 2$ highly mixed nodes and $62\times 0.8871\approx 55$ highly pure nodes. This data is quite unbalanced for its small $\varsigma$.
  • 3.
    For UKfaculty, it has $79\times 0.1139\approx 9$ highly mixed nodes and $79\times 0.6582\approx 52$ highly pure nodes. It is more balanced than the Dolphins and more unbalanced than the Karate club.
  • 4.
    For Political books, it has $92\times 0.0217\approx 2$ highly mixed nodes and $92\times 0.9457\approx 87$ highly pure nodes. It is the most balanced network among all networks because its equilibrium index $\varsigma$ is 0.9771, a value close to 1.
  • 5.
    For Political blogs, it has $1222\times 0.0426\approx 52$ highly mixed nodes and $1222\times 0.8331\approx 1018$ highly pure nodes.
  • 6.
    For Slovene Parliamentary Party, it has zero highly mixed nodes and 7 highly pure nodes. It is quite balanced because 0.9199 is a value close to 1.
  • 7.
    For Gahuku-Gama subtribes, it has 4 highly mixed nodes and 8 highly pure nodes. Its equilibrium index is 0.7159, so it is not that balanced.
  • 8.
    For Karate-club-weighted, the results are similar to that of Karate club, and we omit the analysis here.
  • 9.
    For Train bombing and US Top-500 Airport Network, their mixedness, purity, and equilibrium are similar to that of UKfaculty, and we omit the analysis here.
  • 10.
    For Les Misérables, it has $77\times 0.0130\approx 1$ highly mixed node and $77\times 0.8052\approx 62$ highly pure nodes. It is the most unbalanced network among all networks because the value 0.3335 of its equilibrium index is much smaller than that of other networks.
For visibility, Figure 5 displays communities detected by ScD and we also highlight highly mixed nodes for each data.
Table 4: $\zeta_{\mathrm{mixed}},\zeta_{\mathrm{pure}},\varsigma$ detected by ScD for real data in Table 1. For networks with known $K$, ScD uses the true $K$ to estimate mixed memberships. For networks with unknown $K$, ScD uses $K$ estimated by KScD in Table 2 to estimate mixed memberships.
Dataset $\zeta_{\mathrm{mixed}}$ $\zeta_{\mathrm{pure}}$ $\varsigma$
Karate club 0.0882 0.7353 0.9227
Dolphins 0.0323 0.8871 0.5709
UKfaculty 0.1139 0.6582 0.8041
Political books 0.0217 0.9457 0.9771
Political blogs 0.0426 0.8331 0.8288
Slovene Parliamentary Party 0 0.7000 0.9199
Gahuku-Gama subtribes 0.2500 0.5000 0.7159
Karate-club-weighted 0.0294 0.7941 0.9375
Train bombing 0.1406 0.6250 0.8937
Les Misérables 0.0130 0.8052 0.3335
US Top-500 Airport Network 0.1480 0.6700 0.8981
(a) Karate club
(b) Dolphins
(c) UKfaculty
(d) Political books
(e) Political blogs
(f) Slovene Parliamentary Party
(g) Gahuku-Gama subtribes
(h) Karate-club-weighted
(i) Train bombing
(j) Les Misérables
(k) US Top-500 Airport Network
Figure 5: Communities detected by ScD, where $K$ used here is the same as that of Table 4 for each network. Colors indicate home base communities obtained from $\ell$ and the black square indicates highly mixed nodes, where we let $\ell(i)=\mathrm{arg~{}max}_{k\in[K]}\hat{\Pi}(i,k)$ be the home base community for node $i\in[n]$. For visualization, we do not show node labels and edge weights.
7 Conclusion
Overlapping community detection plays an important role in understanding latent community memberships of complex networks. In this article, we present a general, flexible, and interpretable generative model the Degree-Corrected Mixed Membership Distribution-Free (DCMMDF) model for overlapping weighted networks. DCMMDF can be seen as extensions of several previous models in different ways. In particular, the DCMM model is a sub-model of our DCMMDF. The proposed model can model weighted networks in which nodes belong to multiple communities and edge weights can be generated from any distribution as long as the expectation adjacency matrix has a block structure related to the community memberships.
We use an efficient spectral algorithm to estimate the community memberships of weighted networks generated from DCMMDF. We show that the algorithm is asymptotically consistent under mild conditions for any distribution and we can obtain a theoretical guarantee of consistency for a specific distribution from our main theoretical results immediately by analyzing the variance parameter related to the distribution. We present some examples of how to analyze the algorithm’s performance when the adjacency matrix is generated from different distributions under our DCMMDF.
We also present overlapping weighted modularity to evaluate the quality of overlapping community partition for overlapping weighted networks with positive and negative edge weights. We propose a method to determine the number of communities based on the overlapping weighted modularity. Experiments support our theoretical analysis and show that some classical spectral algorithms also successfully estimate community memberships for weighted networks generated from the proposed model. The advantage and effectiveness of our overlapping weighted modularity are supported by the experiments results that our method successfully estimates the number of communities for weighted networks and it outperforms its competitors.
Our work has several implications: First, ground-truth community memberships of overlapping weighted networks generated from DCMMDF provide a reliable way for evaluating mixed membership community detection approaches. Second, algorithms studied in this paper can broaden our understanding of community structure of weighted networks. Third, our overlapping weighted modularity sheds light on evaluating the quality of overlapping community partition for weighted networks with positive and negative edge weights. And last, our KScD provides a baseline for determining the number of communities for weighted networks.
Future works will be studied from four aspects. First, methods with theoretical guarantees should be designed to infer the number of communities for weighted networks generated from DCMMDF. Second, a theoretical guarantee of Mixed-SCORE under DCMMDF should be developed. Third, given a weighted network, analyzing which distribution that edge weights are coming from is an interesting and challenging task. And last, extending DCMMDF and overlapping weighted modularity from un-directed weighted networks to directed weighted networks is appealing.
CRediT authorship contribution statement
The sole author of the paper has performed all the work.
Declaration of competing interest
The author declared no potential conflicts of interest with respect to the research, authorship, and/or publication of this article.
Data availability
Data will be made available on request.
Appendix A Proof of Theorem 1
Proof.
Let $H_{\hat{U}}=\hat{U}^{\prime}U$. Let $H_{\hat{U}}=U_{H_{\hat{U}}}\Sigma_{H_{\hat{U}}}V^{\prime}_{H_{\hat{U}}}$ be the top $K$ singular value decomposition of $H_{\hat{U}}$. Let $\mathrm{sgn}(H_{\hat{U}})=U_{H_{\hat{U}}}V^{\prime}_{H_{\hat{U}}}$. Under the DCMMDF, the following results are useful for our theoretical analysis.
  • 1.
    $\mathbb{E}[A(i,j)-\Omega(i,j)]=0$ under the DCMMDF.
  • 2.
    $\mathbb{E}[(A(i,j)-\Omega(i,j))^{2}]=\mathrm{Var}(A(i,j))\leq\gamma\theta(i)\theta(j)\leq\gamma\theta^{2}_{\mathrm{max}}$.
  • 3.
    $|A(i,j)-\Omega(i,j)|\leq\tau$ for $i,j\in[n]$.
  • 4.
    Let $\mu=\frac{n\|U\|^{2}_{2\rightarrow\infty}}{K}$ be the incoherence parameter of $\Omega$ defined in Equation (4.23) chen2021spectral . By Lemma H.1 of MaoSVM and Lemma A1 of Criterion , we have $\frac{\theta^{2}_{\mathrm{min}}}{\theta^{2}_{\mathrm{max}}K\lambda_{1}(\Pi^{\prime}\Pi)}\leq\|U\|^{2}_{2\rightarrow\infty}\leq\frac{\theta^{2}_{\mathrm{max}}}{\theta^{2}_{\mathrm{min}}\lambda_{K}(\Pi^{\prime}\Pi)}$, which gives $\frac{\theta^{2}_{\mathrm{min}}n}{\theta^{2}_{\mathrm{max}}K^{2}\lambda_{1}(\Pi^{\prime}\Pi)}\leq\mu\leq\frac{\theta^{2}_{\mathrm{max}}n}{\theta^{2}_{\mathrm{min}}K\lambda_{K}(\Pi^{\prime}\Pi)}$. By Condition 1, we have $\mu=O(1)$.
  • 5.
    Let $c_{b}=\frac{\tau}{\theta_{\mathrm{max}}\sqrt{\gamma n/(\mu\mathrm{log}(n))}}$. By Assumption 1, Condition 1, and the fact that $\mu=O(1)$, we have $c_{b}=O(\frac{\tau\sqrt{\mathrm{log}(n)}}{\theta_{\mathrm{max}}\sqrt{\gamma n}})=O(\sqrt{\frac{\tau^{2}\mathrm{log}(n)}{\gamma\theta^{2}_{\mathrm{max}}n}})=O(\sqrt{\frac{\tau^{2}\mathrm{log}(n)}{\gamma\theta_{\mathrm{max}}\|\theta\|_{1}}})\leq O(1)$.
The above five bullets guarantee that Equations (4.28) and (4.29) of Assumption 4.1. in chen2021spectral hold. Therefore, by Theorem 4.2. chen2021spectral , with probability exceeding $1-O(\frac{1}{n^{5}})$, we have
$\|\hat{U}\mathrm{sgn}(H_{\hat{U}})-U\|_{2\rightarrow\infty}\leq C\frac{\theta_{\mathrm{max}}\kappa(\Omega)\sqrt{\mu\gamma K}+\theta_{\mathrm{max}}\sqrt{\gamma K\mathrm{log}(n)}}{\sigma_{K}(\Omega)},$
provided that $\sigma_{K}(\Omega)\gg\theta_{\mathrm{max}}\sqrt{\gamma n\mathrm{log}(n)}$. Since $\mu=O(1)$ and $K=O(1)$, we have
$\|\hat{U}\mathrm{sgn}(H_{\hat{U}})-U\|_{2\rightarrow\infty}=O(\frac{\theta_{\mathrm{max}}\kappa(\Omega)\sqrt{\gamma}+\theta_{\mathrm{max}}\sqrt{\gamma\mathrm{log}(n)}}{\sigma_{K}(\Omega)}).$
For convenience, set $\varpi=\|\hat{U}\hat{U}^{\prime}-UU^{\prime}\|_{2\rightarrow\infty}$ as the row-wise eigenspace error. Since $\|\hat{U}\hat{U}^{\prime}-UU^{\prime}\|_{2\rightarrow\infty}\leq 2\|U-\hat{U}\mathrm{sgn}(H_{\hat{U}})\|_{2\rightarrow\infty}$, we have
$\varpi=O(\frac{\theta_{\mathrm{max}}\kappa(\Omega)\sqrt{\gamma}+\theta_{\mathrm{max}}\sqrt{\gamma\mathrm{log}(n)}}{\sigma_{K}(\Omega)}).$
Since $\sigma_{K}(\Omega)=\sigma_{K}(\Theta\Pi P\Pi^{\prime}\Theta)\geq\theta^{2}_{\mathrm{min}}\sigma_{K}(\Pi P\Pi^{\prime})\geq\theta^{2}_{\mathrm{min}}\sigma_{K}(P)\lambda_{K}(\Pi^{\prime}\Pi)$ and $\sigma_{1}(\Omega)=\sigma_{1}(\Theta\Pi P\Pi^{\prime}\Theta)\leq\theta^{2}_{\mathrm{max}}\sigma_{1}(\Pi P\Pi^{\prime})\leq\theta^{2}_{\mathrm{max}}\sigma_{1}(P)\lambda_{1}(\Pi^{\prime}\Pi)$ under $DCMMDF_{n}(K,P,\Pi,\Theta,\mathcal{F})$, by Condition 1, we have $\kappa(\Omega)=O(1)$, which gives
$\varpi=O(\frac{\theta_{\mathrm{max}}\kappa(\Omega)\sqrt{\gamma}+\theta_{\mathrm{max}}\sqrt{\gamma\mathrm{log}(n)}}{\theta^{2}_{\mathrm{min}}\sigma_{K}(P)\lambda_{K}(\Pi^{\prime}\Pi)})=O(\frac{\kappa(\Omega)\sqrt{\gamma}+\sqrt{\gamma\mathrm{log}(n)}}{\theta_{\mathrm{min}}\sigma_{K}(P)n})=O(\frac{\sqrt{\gamma\mathrm{log}(n)}}{\theta_{\mathrm{min}}\sigma_{K}(P)n})=O(\frac{\sqrt{\gamma\mathrm{log}(n)}}{\sigma_{K}(P)n\sqrt{\rho}}).$ (8)
Corollary 2 of Criterion is a distribution-free theoretical result, it says that there exists a permutation matrix $\mathcal{P}\in\mathbb{R}^{K\times K}$ such that for $i\in[n]$,
$\|e^{\prime}_{i}(\hat{\Pi}-\Pi\mathcal{P})\|_{1}=O(\varpi\sqrt{n})=O(\frac{\sqrt{\gamma\mathrm{log}(n)}}{\sigma_{K}(P)\sqrt{\rho n}}).$ (9)
Finally, by Condition 1, since $\sigma_{K}(\Omega)\geq\theta^{2}_{\mathrm{min}}\sigma_{K}(P)\lambda_{K}(\Pi^{\prime}\Pi)=O(\sigma_{K}(P)\rho n)$. To make the requirement $\sigma_{K}(\Omega)\gg\theta_{\mathrm{max}}\sqrt{\gamma n\mathrm{log}(n)}=O(\sqrt{\gamma\rho n\mathrm{log}(n)})$ hold, we only need $\sigma_{K}(P)\rho n\gg\sqrt{\gamma\rho n\mathrm{log}(n)}\Leftrightarrow\sigma_{K}(P)\gg\sqrt{\frac{\gamma\mathrm{log}(n)}{\rho n}}$, and this holds naturally because we need the theoretical upper bound of ScD’s error rate provided in Equation (9) go to zero. ∎
6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">